# Haoran Guan

### v1

@@ 2.2.2 Multi-sketching
Moreover, similar to Lemma 5 and Lemma 6, we can derive the following results.
Lemma 8 (Multi-sketching of the matrix)
With similar settings in Lemma 7, for $T\in\mathbb{R}^{m\times n}$, we can have
$\sqrt{1-\epsilon_{s}}\cdot\|T\|_{2}\leq\|\Omega_{2}\Omega_{1}T\|_{2}\leq\sqrt{1+\epsilon_{b}}\cdot\|T\|_{2},$ (12)
$\sqrt{1-\epsilon_{s}}\cdot\|T\|_{F}\leq\|\Omega_{2}\Omega_{1}T\|_{F}\leq\sqrt{1+\epsilon_{b}}\cdot\|T\|_{F},$ (13)
$\sqrt{1-\epsilon_{s}}\cdot\sigma_{min}(T)\leq\sigma_{min}(\Omega_{2}\Omega_{1}T)\leq\|\Omega_{2}\Omega_{1}T\|_{2}\leq\sqrt{1+\epsilon_{b}}\cdot\|T\|_{2},$ (14)
$\frac{\sigma_{min}(\Omega T)}{\sqrt{1+\epsilon_{b}}}\leq\sigma_{min}(T)\leq\|T\|_{2}\leq\frac{\|\Omega T\|_{2}}{\sqrt{1-\epsilon_{s}}}$ (15)
with probability at least $1-p_{f}$. With (15), we can have
$\kappa_{2}(T)\leq\sqrt{\frac{1-\epsilon_{s}}{1+\epsilon_{b}}}\cdot\kappa_{2}(\Omega T)$ (16)
@@
Among all sketching techniques, CountSketch offers low computational and storage costs. However, as indicated by (10), it requires a relatively large sketching size $s$ to maintain the $\epsilon$-embedding properties, which can result in significantly higher FLOPs after HouseholderQR. In contrast, the Gaussian sketch requires a much smaller sketching size than CountSketch, as shown in (9). However, it occupies more space and takes longer to sketch the dense matrix $X$. By applying a Gaussian sketch after a CountSketch, we can achieve low computational and storage costs while leaving a smaller matrix for HouseholderQR, thereby enhancing the algorithm’s efficiency. This is the primary advantage of multi-sketching over traditional single sketching. In this work, we focus on multi-sketching with a Gaussian sketch following a CountSketch. The theoretical analysis steps for other types of sketches are similar and can be explored by the readers.
3 Deterministic LU-Householder CholeskyQR
@@
3.1 Theoretical results of LHC2 and general settings
@@
$PX$ $=$ $LU+E_{LU},$ (17)
$L$ $=$ $WG+E_{HQ},$ (18)
$R$ $=$ $GU-E_{R},$ (19)
$X$ $=$ $QR-E_{QR}.$ (20)
If $X\in\mathbb{R}^{m\times n}$, we make the following assumptions:
$mn\sqrt{n}{\bf u}$$\leq$ $\frac{1}{64},$ (21)
$n^{2}{\bf u}$$\leq$ $\frac{1}{64},$ (22)
$\kappa_{2}(L)\sqrt{mn\sqrt{n}{\bf u}}$$\leq$ $1,$ (23)
$64\kappa_{2}(L)\kappa_{2}(U)n^{2}{\bf u}$$\leq$ 1.1 1.1 . (24)
@@
Theorem 3.1 (Rounding error analysis of LHC2)
If $[Q_{1},R_{2}]=LHC2(X)$, $X\in\mathbb{R}^{m\times n}$, based on our previous settings, we can have
$\|Q_{1}^{\top}Q_{1}-I\|_{F}$$\leq$ $34.85(mn{\bf u}+n(n+1){\bf u}),$ (25)
$\|Q_{1}R_{2}-X\|_{F}$$\leq$ $7.94n^{2}{\bf u}\|X\|_{2}.$ (26)
3.2 Preparations for proving Theorem 3.1
Before proving this theorem, we do some preparations and analysis first. We focus on LHC first. With (20), we can have
$Q^{\top}Q$ $=$ $R^{-\top}(X+E_{QR})^{\top}(X+E_{QR})R^{-1}$ (27)
@@
With (27), we can further decompose $R^{-\top}X^{\top}XR^{-1}$. Based on (17), we can have
$R^{-\top}X^{\top}XR^{-1}$ $=$ $R^{-\top}(LU+E_{LU})^{\top}(LU+E_{LU})R^{-1}$ (28)
@@
We try to transform $R^{-\top}U^{\top}L^{\top}LUR^{-1}$ in (28). With (18) and (19), when $W$ is an orthogonal factor, we can get
$R^{-\top}U^{\top}L^{\top}LUR^{-1}$ $=$ $R^{-\top}U^{\top}(WG+E_{HQ})^{\top}(WG+E_{HQ})UR^{-1}$ (29)
@@
$\|E_{3}\|_{2}$$\leq$ $2\|R^{-1}\|_{2}\|E_{R}\|_{2}+\|R^{-1}\|_{2}^{2}\|E_{R}\|_{2}^{2}.$ (36)
Therefore, with (34)-(36), if we want to estimate (30), we need to estimate $\|R^{-1}\|_{2}$, $\|E_{QR}\|_{2}$, $\|XR^{-1}\|_{2}$, $\|UR^{-1}\|_{2}$, $\|E_{HQ}\|_{2}$, $\|GUR^{-1}\|_{2}$, $\|LUR^{-1}\|_{2}$ and $\|E_{R}\|_{2}$. We do the corresponding estimations below.
Lemma 9 (Estimation of $\|E_{HQ}\|_{2}$)
In (18), we can have
$\|E_{HQ}\|_{2}\leq 1.1mn\sqrt{n}{\bf u}\|L\|_{2}.$ (37)
Proof
According to Lemma 3 and (18), when $W\in\mathbb{R}^{m\times n}$ is orthogonal and $G\in\mathbb{R}^{n\times n}$, we can have
$\|E_{HQj}\|_{2}\leq\gamma_{mn}\|L_{j}\|_{2}.$ (38)
@@
$\leq$ $1.1mn\sqrt{n}{\bf u}\|L\|_{2}.$
(37) is proved.
Lemma 10 (Estimation of $\|E_{R}\|_{2}$)
In (19), we can have
$\|E_{R}\|_{2}\leq 1.13n^{2}{\bf u}\|L\|_{2}\|U\|_{2}.$ (39)
Proof
According to Lemma 2 and (19), we can have
$\|E_{R}\|_{2}\leq 1.1n^{2}{\bf u}\|G\|_{2}\|U\|_{2}.$ (40)
For $\|G\|_{2}$, with (18) and (21), when $W$ is orthogonal, we can have
$\|G\|_{2}$ $=$ $\|WG\|_{2}$ (41)
@@
We put (41) into (40) and we can get
$\|E_{R}\|_{2}$$\leq$ $1.1n^{2}{\bf u}\|U\|_{2}\cdot 1.02\|L\|_{2}$
$\leq$ $1.13n^{2}{\bf u}\|L\|_{2}\|U\|_{2}.$
(39) is proved.
Lemma 11 (Estimation of $\|R^{-1}\|_{2}$)
For $\|R^{-1}\|_{2}$, we can have
$\|R^{-1}\|_{2}\leq\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)}.$ (42)
Proof
Based on Lemma 1 and (19), we can have
$\sigma_{min}(R)$ $=$ $\sigma_{min}(GU-E_{R})$ (43)
@@
For $\sigma_{min}(G)$, with Lemma 1, (18), (21), (23) and (37), we can have
$\sigma_{min}(G)$ $=$ $\sigma_{min}(WG)$ (44)
@@
We put (44) into (43) and combine them with (21)-(24). Therefore, we can get
$\sigma_{min}(R)$ $=$ $0.8\sigma_{min}(L)\sigma_{min}(U)-1,13n^{2}{\bf u}\|L\|_{2}\|U\|_{2}$ (45)
@@
With (45), we can easily receive (42). (42) is proved.
Lemma 12 (Estimation of $\|UR^{-1}\|_{2}$)
For $\|UR^{-1}\|_{2}$, we can have
$\|UR^{-1}\|_{2}\leq\frac{1.29}{\sigma_{min}(L)}.$ (46)
Proof
We multiply both sides of (19) with $R^{-1}$ and we can get
$I+E_{R}R^{-1}=GUR^{-1}.$ (47)
Here, $I\in\mathbb{R}^{n\times n}$ is the identity matrix. We multiply both sides of (47) with $G^{-1}$ and take norms. Thus, we can receive
$\|UR^{-1}\|_{2}=\|G^{-1}\|_{2}(1+\|E_{R}\|_{2}\|R^{-1}\|_{2}).$ (48)
With (44), we can get
$\|G^{-1}\|_{2}\leq\frac{1.25}{\sigma_{min}(L)}.$ (49)
We put (39), (42) and (49) into (48) and with (24), we can get
@@
$\leq$ $\frac{1.25}{\sigma_{min}(L)}\cdot(1+1.13n^{2}{\bf u}\|L\|_{2}\|U\|_{2}\cdot\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)})$
$\leq$ $\frac{1.25}{\sigma_{min}(L)}\cdot(1+1.45n^{2}{\bf u}\kappa_{2}(L)\kappa_{2}(U))$
@@
(46) is proved.
Lemma 13 (Estimation of $\|LUR^{-1}\|_{2}$)
For $\|LUR^{-1}\|_{2}$, we can have
$\|LUR^{-1}\|_{2}\leq 2.$ (50)
Proof
With (18) and (19), when $W$ is orthogonal, we can get
$\|LUR^{-1}\|_{2}^{2}$ $=$ $\|R^{-\top}U^{\top}L^{\top}LUR^{-1}\|_{2}$ (51)
@@
Here, we put (21), (23), (24), (37), (39), (41), (42) and (46) into (51) and we can get
$\|LUR^{-1}\|_{2}^{2}$$\leq$ $(1+\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)}\cdot 1.13n^{2}{\bf u}\|L\|_{2}\|U\|_{2})^{2}$ (52)
$+$ $(\frac{1.29}{\sigma_{min}(L)})^{2}\cdot(1.21m^{2}n^{3}{\bf u}^{2}\|L\|_{2}^{2}+2\cdot 1.02\|L\|_{2}\cdot 1.1mn\sqrt{n}{\bf u}\|L\|_{2})$
@@
With (52), (50) is proved.
Lemma 14 (Estimation of $\|GUR^{-1}\|_{2}$)
For $\|GUR^{-1}\|_{2}$, we can have
$\|GUR^{-1}\|_{2}\leq 1.03.$ (53)
Proof
Based on (47), with (24), (39) and (42), we can get
@@
$\leq$ $1+1.13n^{2}{\bf u}\|L\|_{2}\|U\|_{2}\cdot\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)}$
@@
(53) is proved.
Lemma 15 (Estimation of $\|XR^{-1}\|_{2}$)
For $\|XR^{-1}\|_{2}$, we can have
$\|XR^{-1}\|_{2}\leq 2.02.$ (54)
Proof
With (17), we can have
$\|XR^{-1}\|_{2}^{2}$ $=$ $\|R^{-\top}X^{\top}XR^{-1}\|_{2}$ (55)
@@
To estimate $\|E_{LU}\|_{2}$, with Lemma 2 and (17), we can have
$\|E_{LU}\|_{2}\leq 1.1n^{2}{\bf u}\|L\|_{2}\|U\|_{2}.$ (56)
We put (42), (50) and (56) into (55) and with (24), we can have
$\|XR^{-1}\|_{2}^{2}$$\leq$ $\|LUR^{-1}\|_{2}^{2}+2\|E_{LU}\|_{2}\|R^{-1}\|_{2}\|LUR^{-1}\|_{2}+\|E_{LU}\|_{2}^{2}\|R^{-1}\|_{2}^{2}$ (57)
$\leq$ $3.97+2\cdot 1.1n^{2}{\bf u}\|L\|_{2}\|U\|_{2}\cdot\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)}\cdot 2$
$+$ $(1.1n^{2}{\bf u}\|L\|_{2}\|U\|_{2})^{2}\cdot(\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)})^{2}$
@@
With (57), we can have (54). (54) is proved.
Lemma 16 (Estimation of $\|E_{QR}\|_{2}$)
For $\|E_{QR}\|_{2}$, we can have
$\|E_{QR}\|_{2}\leq 2.33n^{2}{\bf u}\|R\|_{2}.$ (58)
Proof
With Lemma2, (20), we can have
$\|E_{QR}\|_{2}\leq 1.1n^{2}\|Q\|_{2}\|R\|_{2}.$ (59)
For $\|Q\|_{2}$, based on Lemma 2 and (20), we can have
$\|Q\|_{2}$$\leq$ $\|XR^{-1}\|_{2}+\|E_{QR}R^{-1}\|_{2}$ (60)
$\leq$ $\|XR^{-1}\|_{2}+1.1n^{2}{\bf u}\|Q\|_{2}\|R\|_{2}\cdot\|R^{-1}\|_{2}$
$\leq$ $\|XR^{-1}\|_{2}+1.1n^{2}{\bf u}\kappa_{2}(R)\|Q\|_{2}.$
For $\kappa_{2}(R)$, we need to estimate $\|R\|_{2}$ first. With (19), (22), (39) and (41), we can get
$\|R\|_{2}$$\leq$ $\|G\|_{2}\|U\|_{2}+\|E_{R}\|_{2}$ (61)
$\leq$ $1.02\|L\|_{2}\|U\|_{2}+1.13n^{2}{\bf u}\|L\|_{2}\|U\|_{2}$
@@
We combine (42) with (61) and we can estimate $\kappa_{2}(R)$ as
$\kappa_{2}(R)$ $=$ $\|R\|_{2}\|R^{-1}\|_{2}$ (62)
@@
We put (54) and (62) into (63) and with (24), we can have
$\|Q\|_{2}$$\leq$ $\frac{\|XR^{-1}\|_{2}}{1-1.1n^{2}{\bf u}\kappa_{2}(R)}$ (63)
@@
We put (62) and (63) into (59) and we can have (58). (58) is proved.
3.3 Proof of Theorem 3.1
With all the lemmas above, we start to prove Theorem 3.1. Some steps and ideas are similar to that of LUChol but there are also many differences.
Proof
Our proof of Theorem 3.1 has two parts, orthogonality and residual.
3.3.1 Orthogonality
We estimate the orthogonality of LHC2 first. According to (30), we estimate $\|E_{1}\|_{2}$, $\|E_{2}\|_{2}$ and $\|E_{3}\|_{2}$ seperately.
For $\|E_{1}\|_{2}$, we put (42), (54) and (58) into (34). With (24) and (61), we can get
$\|E_{1}\|_{2}$$\leq$ $2\|R^{-1}\|_{2}\|E_{QR}\|_{2}\|XR^{-1}\|_{2}+\|R^{-1}\|_{2}^{2}\|E_{QR}\|_{2}^{2}$ (64)
$\leq$ $2\cdot\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)}\cdot 2.33n^{2}{\bf u}\|R\|_{2}\cdot 2.02$
$+$ $(\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)})^{2}\cdot(2.33n^{2}{\bf u}\|R\|_{2})^{2}$
$\leq$ $2\cdot 1.28\cdot 1.04\cdot 2.33n^{2}{\bf u}\kappa_{2}(L)\kappa_{2}(U)\cdot 2.02$
$+$ $(1.28\cdot 2.33\cdot 1.04)^{2}\cdot(n^{2}{\bf u}\kappa_{2}(L)\kappa_{2}(U))^{2}$

### v2

@@ 2.2.2. Multi-sketching
Moreover, similar to Lemma 2.6 and Lemma 2.7, we can derive the following results.
Lemma 2.10 (Multi-sketching of the matrix).
With similar settings in Lemma 2.9, for $T\in\mathbb{R}^{m\times n}$, we have
(2.12) $\sqrt{1-\epsilon_{s}}\cdot\|T\|_{2}\leq\|\Omega_{2}\Omega_{1}T\|_{2}\leq\sqrt{1+\epsilon_{b}}\cdot\|T\|_{2},$
(2.13) $\sqrt{1-\epsilon_{s}}\cdot\|T\|_{F}\leq\|\Omega_{2}\Omega_{1}T\|_{F}\leq\sqrt{1+\epsilon_{b}}\cdot\|T\|_{F},$
(2.14) $\sqrt{1-\epsilon_{s}}\cdot\sigma_{min}(T)\leq\sigma_{min}(\Omega_{2}\Omega_{1}T)\leq\|\Omega_{2}\Omega_{1}T\|_{2}\leq\sqrt{1+\epsilon_{b}}\cdot\|T\|_{2},$
(2.15) $\frac{\sigma_{min}(\Omega T)}{\sqrt{1+\epsilon_{b}}}\leq\sigma_{min}(T)\leq\|T\|_{2}\leq\frac{\|\Omega T\|_{2}}{\sqrt{1-\epsilon_{s}}}$
with probability at least $1-p_{f}$. With (2.15), we have
(2.16) $\kappa_{2}(T)\leq\sqrt{\frac{1-\epsilon_{s}}{1+\epsilon_{b}}}\cdot\kappa_{2}(\Omega T)$
@@
Among all sketching techniques, CountSketch offers low computational and storage costs. However, as indicated by (2.10), it requires a relatively large sketching size $s$ to maintain the $\epsilon$-embedding properties, which can result in significantly higher FLOPs after HouseholderQR. In contrast, the Gaussian sketch requires a much smaller sketching size than CountSketch, as shown in (2.9). However, it occupies more space and takes longer to sketch the dense matrix $X$. By applying a Gaussian sketch after a CountSketch, we can achieve low computational and storage costs while leaving a smaller matrix for HouseholderQR, thereby enhancing the algorithm’s efficiency. This is the primary advantage of multi-sketching over traditional single sketching. In this work, we focus on multi-sketching with a Gaussian sketch following a CountSketch. The theoretical analysis steps for other types of sketches are similar and can be explored by the readers.
3. Deterministic LU-Householder CholeskyQR
@@
3.1. Theoretical results of LHC2 and general settings
@@
(3.1) $PX$ $=$ $LU+E_{LU},$
(3.2) $L$ $=$ $WG+E_{HQ},$
(3.3) $R$ $=$ $GU-E_{R},$
(3.4) $X$ $=$ $QR-E_{QR}.$
If $X\in\mathbb{R}^{m\times n}$, we make the following settings:
(3.5) $mn\sqrt{n}{\bf u}$$\leq$ $\frac{1}{64},$
(3.6) $n^{2}{\bf u}$$\leq$ $\frac{1}{64},$
(3.7) $\kappa_{2}(L)\sqrt{mn\sqrt{n}{\bf u}}$$\leq$ $1,$
(3.8) $64\kappa_{2}(L)\kappa_{2}(U)n^{2}{\bf u}$$\leq$ 1.1 1.1 .
@@
Theorem 3.1 (Rounding error analysis of LHC2).
For $X\in\mathbb{R}^{m\times n}$ and $[Q_{1},R_{2}]=\mbox{LHC2}(X)$, when (3.5)-(3.8) are satisfied, we have
(3.9) $\|Q_{1}^{\top}Q_{1}-I\|_{F}$$\leq$ $34.85(mn{\bf u}+n(n+1){\bf u}),$
(3.10) $\|Q_{1}R_{2}-X\|_{F}$$\leq$ $7.94n^{2}{\bf u}\|X\|_{2}.$
3.2. Preparations for proving Theorem 3.1
Before proving this theorem, we do some preparations and analysis first. We focus on LHC first. With (3.4), we can have
(3.11) $Q^{\top}Q$ $=$ $R^{-\top}(X+E_{QR})^{\top}(X+E_{QR})R^{-1}$
@@
With (3.11), we can further decompose $R^{-\top}X^{\top}XR^{-1}$. Based on (3.1), we can have
(3.12) $R^{-\top}X^{\top}XR^{-1}$ $=$ $R^{-\top}(LU+E_{LU})^{\top}(LU+E_{LU})R^{-1}$
@@
We try to transform $R^{-\top}U^{\top}L^{\top}LUR^{-1}$ in (3.12). With (3.2) and (3.3). When $W$ is orthogonal, we can get
(3.13) $R^{-\top}U^{\top}L^{\top}LUR^{-1}$ $=$ $R^{-\top}U^{\top}(WG+E_{HQ})^{\top}(WG+E_{HQ})UR^{-1}$
@@
(3.20) $\|E_{3}\|_{2}$$\leq$ $2\|R^{-1}\|_{2}\|E_{R}\|_{2}+\|R^{-1}\|_{2}^{2}\|E_{R}\|_{2}^{2}.$
We can see from (3.18)-(3.20) that if we want to estimate $\|Q^{\top}Q-I\|_{2}$ in (3.14), we need to estimate $\|R^{-1}\|_{2}$, $\|E_{QR}\|_{2}$, $\|XR^{-1}\|_{2}$, $\|UR^{-1}\|_{2}$, $\|E_{HQ}\|_{2}$, $\|GUR^{-1}\|_{2}$, $\|LUR^{-1}\|_{2}$ and $\|E_{R}\|_{2}$. We provide the corresponding estimations below.
Lemma 3.2 (Estimation of $\|E_{HQ}\|_{2}$).
For $\|E_{HQ}\|_{2}$, we have
(3.21) $\|E_{HQ}\|_{2}\leq 1.1mn\sqrt{n}{\bf u}\cdot\|L\|_{2}.$
Proof.
According to Lemma 2.3 and (3.2), when $W\in\mathbb{R}^{m\times n}$ is orthogonal, we can have
(3.22) $\|E_{HQj}\|_{2}\leq\gamma_{mn}\|L_{j}\|_{2}.$
@@
$\leq$ $1.1mn\sqrt{n}{\bf u}\cdot\|L\|_{2}.$
(3.21) is proved. ∎
Lemma 3.3 (Estimation of $\|E_{R}\|_{2}$).
For $\|E_{R}\|_{2}$, we have
(3.23) $\|E_{R}\|_{2}\leq 1.13n^{2}{\bf u}\cdot\|L\|_{2}\|U\|_{2}.$
Proof.
According to Lemma 2.2 and (3.3), we can have
(3.24) $\|E_{R}\|_{2}$$\leq$ $\gamma_{n}\cdot\|G\|_{F}\|U\|_{F}$
$\leq$ $\gamma_{n}\cdot n\|G\|_{2}\|U\|_{2}$
$\leq$ $1.1n^{2}{\bf u}\cdot\|G\|_{2}\|U\|_{2}.$
For $\|G\|_{2}$, with (3.2) and (3.5), when $W$ is orthogonal, we can have
(3.25) $\|G\|_{2}$ $=$ $\|WG\|_{2}$
@@
We put (3.25) into (3.24) and we can get
$\|E_{R}\|_{2}$$\leq$ $1.1n^{2}{\bf u}\cdot\|U\|_{2}\cdot 1.02\|L\|_{2}$
$\leq$ $1.13n^{2}{\bf u}\cdot\|L\|_{2}\|U\|_{2}.$
(3.23) is proved. ∎
Lemma 3.4 (Estimation of $\|R^{-1}\|_{2}$).
For $\|R^{-1}\|_{2}$, we have
(3.26) $\|R^{-1}\|_{2}\leq\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)}.$
Proof.
Based on Lemma 2.1 and (3.3), we can have
(3.27) $\sigma_{min}(R)$ $=$ $\sigma_{min}(GU-E_{R})$
@@
For $\sigma_{min}(G)$, with Lemma 2.1, (3.2), (3.5), (3.7) and (3.21), we can have
(3.28) $\sigma_{min}(G)$ $=$ $\sigma_{min}(WG)$
@@
We put (3.28) into (3.27) and with (3.5)-(3.8), we can get
(3.29) $\sigma_{min}(R)$ $=$ $0.8\sigma_{min}(L)\sigma_{min}(U)-1,13n^{2}{\bf u}\cdot\|L\|_{2}\|U\|_{2}$
@@
With (3.29), we can easily get (3.26). (3.26) is proved. ∎
Lemma 3.5 (Estimation of $\|UR^{-1}\|_{2}$).
For $\|UR^{-1}\|_{2}$, we have
(3.30) $\|UR^{-1}\|_{2}\leq\frac{1.29}{\sigma_{min}(L)}.$
Proof.
We multiply both sides of (3.3) with $R^{-1}$ and we can get
(3.31) $I+E_{R}R^{-1}=GUR^{-1}.$
Here, $I\in\mathbb{R}^{n\times n}$ is the identity matrix. We multiply both sides of (3.31) with $G^{-1}$ and we can have
(3.32) $\|UR^{-1}\|_{2}=\|G^{-1}\|_{2}(1+\|E_{R}\|_{2}\|R^{-1}\|_{2}).$
With (3.28), we can get
(3.33) $\|G^{-1}\|_{2}\leq\frac{1.25}{\sigma_{min}(L)}.$
We put (3.23), (3.26) and (3.33) into (3.32) and with (3.8), we can get
@@
$\leq$ $\frac{1.25}{\sigma_{min}(L)}\cdot(1+1.13n^{2}{\bf u}\cdot\|L\|_{2}\|U\|_{2}\cdot\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)})$
$\leq$ $\frac{1.25}{\sigma_{min}(L)}\cdot(1+1.45n^{2}{\bf u}\cdot\kappa_{2}(L)\kappa_{2}(U))$
@@
(3.30) is proved. ∎
Lemma 3.6 (Estimation of $\|LUR^{-1}\|_{2}$).
For $\|LUR^{-1}\|_{2}$, we have
(3.34) $\|LUR^{-1}\|_{2}\leq 2.$
Proof.
With (3.2) and (3.3), when $W$ is orthogonal, we can get
(3.35) $\|LUR^{-1}\|_{2}^{2}$ $=$ $\|R^{-\top}U^{\top}L^{\top}LUR^{-1}\|_{2}$
@@
Here, we put (3.5), (3.7), (3.8), (3.21), (3.23), (3.25), (3.26) and (3.30) into (3.35) and we can get
(3.36) $\|LUR^{-1}\|_{2}^{2}$$\leq$ $(1+\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)}\cdot 1.13n^{2}{\bf u}\cdot\|L\|_{2}\|U\|_{2})^{2}$
$+$ $(\frac{1.29}{\sigma_{min}(L)})^{2}\cdot(1.21m^{2}n^{3}{\bf u}^{2}\cdot\|L\|_{2}^{2}+2\cdot 1.02\|L\|_{2}\cdot 1.1mn\sqrt{n}{\bf u}\cdot\|L\|_{2})$
@@
With (3.36), (3.34) is proved. ∎
Lemma 3.7 (Estimation of $\|GUR^{-1}\|_{2}$).
For $\|GUR^{-1}\|_{2}$, we have
(3.37) $\|GUR^{-1}\|_{2}\leq 1.03.$
Proof.
Based on (3.31), with (3.8), (3.23) and (3.26), we can get
@@
$\leq$ $1+1.13n^{2}{\bf u}\cdot\|L\|_{2}\|U\|_{2}\cdot\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)}$
@@
(3.37) is proved. ∎
Lemma 3.8 (Estimation of $\|XR^{-1}\|_{2}$).
For $\|XR^{-1}\|_{2}$, we have
(3.38) $\|XR^{-1}\|_{2}\leq 2.02.$
Proof.
With (3.1), we can have
(3.39) $\|XR^{-1}\|_{2}^{2}$ $=$ $\|R^{-\top}X^{\top}XR^{-1}\|_{2}$
@@
To estimate $\|E_{LU}\|_{2}$, with Lemma 2.2 and (3.1), we can have
(3.40) $\|E_{LU}\|_{2}$$\leq$ $\gamma_{n}\cdot\|L\|_{F}\|U\|_{F}$
$\leq$ $\gamma_{n}\cdot n\|L\|_{2}\|U\|_{2}$
$\leq$ $1.1n^{2}{\bf u}\cdot\|L\|_{2}\|U\|_{2}.$
We put (3.26), (3.34) and (3.40) into (3.39) and with (3.8), we can have
(3.41) $\|XR^{-1}\|_{2}^{2}$$\leq$ $\|LUR^{-1}\|_{2}^{2}+2\|E_{LU}\|_{2}\|R^{-1}\|_{2}\|LUR^{-1}\|_{2}+\|E_{LU}\|_{2}^{2}\|R^{-1}\|_{2}^{2}$
$\leq$ $3.97+2\cdot 1.1n^{2}{\bf u}\cdot\|L\|_{2}\|U\|_{2}\cdot\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)}\cdot 2$
$+$ $(1.1n^{2}{\bf u}\cdot\|L\|_{2}\|U\|_{2})^{2}\cdot(\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)})^{2}$
@@
With (3.41), we can have (3.38). (3.38) is proved. ∎
Lemma 3.9 (Estimation of $\|E_{QR}\|_{2}$).
For $\|E_{QR}\|_{2}$, we have
(3.42) $\|E_{QR}\|_{2}\leq 2.33n^{2}{\bf u}\cdot\|R\|_{2}.$
Proof.
With Lemma2.2, (3.4), we can have
(3.43) $\|E_{QR}\|_{2}$$\leq$ $\gamma_{n}\cdot\|Q\|_{F}\|R\|_{F}$
$\leq$ $\gamma_{n}\cdot n\|Q\|_{2}\|R\|_{2}$
$\leq$ $1.1n^{2}{\bf u}\cdot\|Q\|_{2}\|R\|_{2}.$
For $\|Q\|_{2}$, based on Lemma 2.2 and (3.4), we can have
(3.44) $\|Q\|_{2}$$\leq$ $\|XR^{-1}\|_{2}+\|E_{QR}R^{-1}\|_{2}$
$\leq$ $\|XR^{-1}\|_{2}+1.1n^{2}{\bf u}\cdot\|Q\|_{2}\|R\|_{2}\cdot\|R^{-1}\|_{2}$
$\leq$ $\|XR^{-1}\|_{2}+1.1n^{2}{\bf u}\cdot\kappa_{2}(R)\|Q\|_{2}.$
For $\kappa_{2}(R)$, we need to estimate $\|R\|_{2}$ first. With (3.3), (3.6), (3.23) and (3.25), we can get
(3.45) $\|R\|_{2}$$\leq$ $\|G\|_{2}\|U\|_{2}+\|E_{R}\|_{2}$
$\leq$ $1.02\|L\|_{2}\|U\|_{2}+1.13n^{2}{\bf u}\cdot\|L\|_{2}\|U\|_{2}$
@@
We combine (3.26) with (3.45) and we can estimate $\kappa_{2}(R)$ as
(3.46) $\kappa_{2}(R)$ $=$ $\|R\|_{2}\|R^{-1}\|_{2}$
@@
We put (3.38) and (3.46) into (3.47) and based on (3.6) and (3.8), we can have
(3.47) $\|Q\|_{2}$$\leq$ $\frac{\|XR^{-1}\|_{2}}{1-1.1n^{2}{\bf u}\cdot\kappa_{2}(R)}$
@@
We put (3.47) into (3.43) and we can have (3.42). (3.42) is proved. ∎
3.3. Proof of Theorem 3.1
With all the lemmas above, we start to prove Theorem 3.1. Some steps and ideas are similar to those in [34].
Proof.
Our proof of Theorem 3.1 consists of two parts, estimating the orthogonality and estimating the residual.
3.3.1. Orthogonality
We estimate the orthogonality of LHC2 first. According to (3.14), we estimate $\|E_{1}\|_{2}$, $\|E_{2}\|_{2}$ and $\|E_{3}\|_{2}$ separately. For $\|E_{1}\|_{2}$, we put (3.26), (3.38) and (3.42) into (3.18) and with (3.8) and (3.45), we can get
(3.48) $\|E_{1}\|_{2}$$\leq$ $2\|R^{-1}\|_{2}\|E_{QR}\|_{2}\|XR^{-1}\|_{2}+\|R^{-1}\|_{2}^{2}\|E_{QR}\|_{2}^{2}$
$\leq$ $2\cdot\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)}\cdot 2.33n^{2}{\bf u}\cdot\|R\|_{2}\cdot 2.02$
$+$ $(\frac{1.28}{\sigma_{min}(L)\sigma_{min}(U)})^{2}\cdot(2.33n^{2}{\bf u}\cdot\|R\|_{2})^{2}$
$\leq$ $2\cdot 1.28\cdot 1.04\cdot 2.33n^{2}{\bf u}\cdot\kappa_{2}(L)\kappa_{2}(U)\cdot 2.02$
$+$ $(1.28\cdot 2.33\cdot 1.04)^{2}\cdot(n^{2}{\bf u}\cdot\kappa_{2}(L)\kappa_{2}(U))^{2}$